please explain each step
9) Find is, i2, i3, and is 3A v2 102 40 12 § 6A 40127 0 2 A
Answer #1

GA I 1 A www you to yol for Solving in, iz, i3, iy we have to use Super position Theorem Superposition Theorem if a number ofNow by Current division rule 1 10 - XGA 10+20 = 1/2 X 6A = QA 12 = 20 20+10 GA = 2 x 6A = UA here is, and it are zero because90 - X3A = I x3A = 1.5A qoto yo - X3 A = 3 X 3A = 1.5A 40+40 Now lets take an Current Source alone 3 is 14 OJE ww LO vons soUsing the principle of Superponition theorem. = 2A+ 1A + 0 = 3A 12 = 4A + 2A + 0 = 6A iz = of 1. 5A + (A = 2.5A 1y = 0ut 1.5A